OpenStudy (anonymous):

Jim's sock drawer has two pairs of black socks, three pairs of white, one pair of green, and two pairs of gray. One evening he randomly picks three socks assuming that he will get a matched pair, but instead turns up with 1 black, 1 white, and 1 gray sock. Without putting back the socks he picked, he picks another sock randomly. What is the probability that he will have a complete pair?

OpenStudy (ankit042):

b = 2*2 =4 w = 3*2 =6 green = 1*2 =2 grey = 2*2 =4

OpenStudy (ankit042):

Can you try the next steps?? Hint now as he has removed a few socks the count will change for respective pairs

OpenStudy (anonymous):

so there is 16 socks total but is left with 13 because he takes out 3?

OpenStudy (ankit042):

Yes!! so think of the cases by which he can make a complete pair

OpenStudy (anonymous):

so 3 black 5 white 3 grey

OpenStudy (anonymous):

so 11/13?

OpenStudy (ankit042):

Yes!! see you know this stuff, don't need my help

OpenStudy (anonymous):

the words cunfuzzle me /.\

OpenStudy (anonymous):

thank you!!!!!!!!!!!!!!!!!!!!!!!!